### v1

@@ 3.1. Pencils of quadrics
For any $\Gamma$-invariant Riemannian metric $g$ on $\mathbb{S}\mathcal{E}$, the map $s$ from the proof of Proposition 4.5 is a quasi-isometric embedding. Indeed let $\delta$ be the maximum of $d_{g}(x,x^{\prime})$ or $d_{H}(s(x),s(x^{\prime}))$ for any $x,x^{\prime}$ in the same fiber of $p:\mathbb{S}\mathcal{E}\to\widetilde{N}$. For every $x,y\in\mathbb{S}\mathcal{E}$ one can find $x^{\prime},y^{\prime}$ in the same fibers respectively as $x,y$ and in the same flow line for $\Phi$. Hence for the constants $C,D$ from Proposition 4.5:
$\frac{1}{C}d_{g}(x,y)-D-(\frac{2}{C}+2)\delta\leq d_{H}(s(x),s(y))\leq Cd_{g}(x,y)+D+(2C+2)\delta.$
@@ 4.2. Existence of fitting flows
Recall that $p:\mathcal{E}\to\operatorname{Gr}_{d}(\mathcal{Q})$ is the bundle map and $\pi:\mathcal{E}\to\mathcal{Q}$ is the tautological map.
@@
This lemma is an inverse of point $(iii)$ of Proposition 3.10: here we fix an element $q\in u(x)$ whereas before we were fixing a $v\in TM$. In order to construct this inverse we will use the fact that a continuous odd map between spheres of equal dimensions is surjective.
@@
Finally we will make this construction so that in addition $\mathrm{d}p\left(\psi(v)\right)=v$ for $v\in T_{x}M\setminus\{0\}$, but this property will be only used during the construction.
@@
${Tu^{*}\mathcal{E}}$${T\mathcal{Q}}$${TM}$${u^{*}\mathcal{E}}$${\mathcal{Q}}$${M}$$\scriptstyle{\mathrm{d}m_{\lambda}}$$\scriptstyle{\mathrm{d}\pi}$$\scriptstyle{\mathrm{d}p}$$\scriptstyle{\psi}$$\scriptstyle{\phi}$$\scriptstyle{m_{\lambda}}$$\scriptstyle{\pi}$$\scriptstyle{p}$
If we can construct such continuous maps, the fact that $\overline{\phi}$ is an odd map between two spheres of the same dimension implies that it is homotopically non-trivial and therefore surjective. In particular for all $[q]\in\mathbb{S}\mathcal{E}_{u(x)}\simeq\mathbb{S}u(x)$ there exist a $v\in T_{x}M$ such that $\phi(v)=q$. The element $\mathrm{w}=\psi(v)$ then satisfies the required conditions, so this finishes the proof.
@@
Since $u$ is a fitting immersion, the point $(iii)$ of Proposition 3.10 implies that given $v_{0}\in T_{x}M$ one can find $\phi(v_{0})\in u^{*}\mathcal{E}_{x}$ an $\psi(v_{0})\in T_{\phi(v_{0})}u^{*}\mathcal{E}$ such that $\mathrm{d}\pi(\psi(v_{0}))$ is positive and $\mathrm{d}p\left(\psi(v_{0})\right)=v_{0}$.
Note that the condition that $\mathrm{d}\pi(\psi(v))$ is positive is an open condition and the condition that $\mathrm{d}p\left(\psi(v)\right)=v$ requires that $\psi$ is a section of an affine sub-bundle. Hence for every $v\in T_{x}M\setminus\{0\}$ we can find a small neighborhood $\mathcal{S}$ in a sphere in $T_{x}M$ containing $v_{0}$ on which we can define $\phi$ and a lift $\psi$ such that $\mathrm{d}\pi(\psi(v))$ is positive and $\mathrm{d}p\left(\psi(v)\right)=v$ for all $v\in\mathcal{S}$. We take $\mathcal{S}$ small enough so that it does not contain any antipodal pair of points.
We define $U$ to be the set of non-zero elements $\lambda w$ for all $\lambda\in\mathbb{R}$ and $w\in\mathcal{S}$, and we extend $\phi$ and $\psi$ to $U$ in a homogeneous way. We define $\phi$ on $U$ so that $\phi_{i}(\lambda v)=\lambda\phi_{i}(v)$ for all $\lambda\in\mathbb{R}$ non-zero and $v\in\mathcal{S}$. We set $\psi(\lambda v)=\lambda(\mathrm{d}m_{\lambda})\psi_{i}(v)$ for $\lambda\in\mathbb{R}$, where $m_{\lambda}$ is the multiplication by $\lambda$ on $u^{*}\mathcal{E}_{x}$. Note that $\mathrm{d}\pi\left(\psi_{i}(\lambda v)\right)=\lambda^{2}\mathrm{d}\pi\left(\psi_{i}(v)\right)$ is positive and $\mathrm{d}p\left(\psi_{i}(\lambda v)\right)=\lambda v$ for all $\lambda v\in U$. Indeed $\pi\circ m_{\lambda}=\lambda\pi$ so $\mathrm{d}\pi\circ\mathrm{d}m_{\lambda}=\lambda\mathrm{d}\pi$ and $p\circ m_{\lambda}=p$ so $\mathrm{d}p\circ\mathrm{d}m_{\lambda}=\mathrm{d}p$.
We therefore can construct an open cover $\{U_{i}\}_{i\in I}$ of $T_{x}M$ and continuous maps $\phi_{i}:U_{i}\to\mathbb{S}u^{*}\mathcal{E}\subset u^{*}\mathcal{E}$, and lifts $\psi_{i}:U_{i}\to Tu^{*}\mathcal{E}$ such that $\mathrm{d}\pi(\psi_{i}(v))$ is positive and $\mathrm{d}p\left(\psi_{i}(v)\right)=v$ for any $i\in I$ and $v\in U_{i}$. The $U_{i}$ can be assumed invariant by scalar multiplication, and $\phi$ and $\psi$ satisfy the aforementioned homogeneity conditions.
@@
Let us check that $\phi(v)$ is always non-zero. This is where we use that $\mathrm{d}p\left(\psi_{i}(v)\right)=v$ for $i\in I$, and we also use that the fitting immersion $u$ is defining a smooth fibration of the cone $S^{2}V^{>0}$. Let $\gamma:\mathbb{R}\to M$ be a curve such that $\gamma(0)=x$ and $\gamma^{\prime}(0)=v$. For all $t\in\mathbb{R}$ the intersection $u^{\circ}(\gamma(t))\cap S^{2}V^{>0}$ is a non-empty convex set, so we can construct a section $s:t\in\mathbb{R}\mapsto S^{2}V^{>0}$ of the fibration, i.e. such that for all $t\in\mathbb{R}$, $s(t)\in u^{\circ}(\gamma(t))$.
Let us fix $i\in I$. Let $q:\mathbb{R}\to\mathbb{S}u^{*}\mathcal{E}$ be such that $q^{\prime}(0)=\psi(v)$, which implies that $p\circ q(t)$ is equal to $\gamma(t)$ at the first order around $t=0$ since $\gamma^{\prime}(0)=\mathrm{d}p\left(\psi(v)\right)=v$. Since $s(t)\in u^{\circ}(\gamma(t))$, one has $\pi(q(t))\cdot s(t)=0$ at the first order around $t=0$. Taking the first derivative at $t=0$ of this equation we get:
@@
$\mathrm{d}p(\psi(v))=\sum_{i\in I}\chi_{i}(v)\mathrm{d}p(\psi_{i}(v))=\left(\sum_{i\in I}\chi_{i}(v)\right)v=v.$
@@
To prove Proposition 4.8 we use the directions $\mathrm{w}$ from Lemma 4.9 and we glue these vectors into a vector field using a partition of the unit. We construct the vector field $W:\mathbb{S}u^{*}\mathcal{E}\to T\mathbb{S}u^{*}\mathcal{E}$ in a similar manner as $\psi:\mathbb{S}TM\to T\mathbb{S}u^{*}\mathcal{E}$. Morally "$W=\psi\circ\phi^{-1}$", but the map $\phi$ constructed previously is not a priori bijective.
@@
We construct a continuous vector field $\mathrm{W}$ over $u^{*}\mathcal{E}$, except the zero section, such that for all $x\in M$ and non zero $q\in u(x)$, $W_{\lambda q}=\mathrm{d}(m_{\lambda})W_{q}$ for $\lambda\in\mathbb{R}^{>0}$, where $m_{\lambda}$ is the multiplication by $\lambda$ on $\mathcal{E}$. Such a vector field defines a vector field $\overline{W}$ on $\mathbb{S}u^{*}\mathcal{E}$. We require moreover that $\mathrm{d}\pi(W)$ is always positive.
Given any non-zero $q_{0}\in u^{*}\mathcal{E}$, Lemma 4.9 provides the existence of some $\mathrm{w}\in T_{(u(x),q)}\mathcal{E}$ such that $\mathrm{d}p(\mathrm{w})\in\mathrm{d}u(T_{x}M)$ and $\mathrm{d}\pi(\mathrm{w})$ is positive. The first properties implies that $\mathrm{w}$ defines a vector in $T_{x,q_{0}}u^{*}\mathcal{E}$. For each such $q_{0}$ one can find a neighborhood $U$ of $q$ in $u^{*}\mathcal{E}$ that is invariant by the $\mathbb{R}^{>0}$-action, and on which one can define a map $W$ satisfiying the required properties.
Using a partition of the unit as in Lemma 4.9 we construct the desired vector field on $u^{*}\mathcal{E}$. This vector field generates a flow that is a fitting flow.
@@ 5.1. The Anosov property
Definition 5.1.
Let $q_{1},q_{2}\in\mathbb{S}\mathcal{Q}^{\operatorname{mix}}$ be such that for some choice of representatives, the difference $q_{2}-q_{1}$ is positive. We define the cross-ratio distance $\text{cr}([q_{2}],[q_{1}])$ between $[q_{2}]$ and $[q_{1}]$ as the minimum of $[\ell_{1},\ell_{2},\ell_{2}^{\prime},\ell_{1}^{\prime}]\in[1,\infty]$ for every quadruple $\ell_{1},\ell_{2},\ell_{2}^{\prime},\ell_{1}^{\prime}\in\mathbb{P}(V)$ of points in this order on a projective line $L$ such that the zeros of $q_{1}$ on $L$ with multiplicity are $\ell_{1}$ and $\ell_{1}^{\prime}$ and the zeros of $q_{2}$ on $L$ are $\ell_{2}$ and $\ell_{2}^{\prime}$.
If we fix an affine chart of $L$ so that $\ell_{1},\ell_{2},\ell_{2}^{\prime},\ell_{1}^{\prime}$ correspond to the real numbers $x_{2}^{\prime}<x_{1}^{\prime}\leq x_{1}<x_{2}$, the cross ratio $[\ell_{1},\ell_{2},\ell_{2}^{\prime},\ell_{1}^{\prime}]$ is defined as:
$[\ell_{1},\ell_{2},\ell_{2}^{\prime},\ell_{1}^{\prime}]=\frac{x_{2}-x_{1}^{\prime}}{x_{2}-x_{2}^{\prime}}\times\frac{x_{1}-x_{2}^{\prime}}{x_{1}-x_{1}^{\prime}}.$
Note that this quantity is greater than $1$, and is infinite if and only if $\ell_{1}=\ell_{1}^{\prime}$ or $\ell_{2}=\ell_{2}^{\prime}$. However if $q_{1}$ and $q_{2}$ are not mixed, there exist a projective line $L$ between a point on which $q_{2}$ is positive and a point on which $q_{1}$ is negative. Therefore on this line $\ell_{1}\neq\ell_{1}^{\prime}$ and $\ell_{2}\neq\ell_{2}^{\prime}$, so $\text{cr}([q_{2}],[q_{1}])$ is finite.
@@
Let $[q_{1}],[q_{2}],[q_{3}]\in\mathbb{S}\mathcal{Q}^{\operatorname{mix}}$ be such that $q_{3}-q_{2}$ and $q_{2}-q_{1}$ are positive. Then :
@@
Pick any projective line that crosses the zeroes of the quadric $q_{1}$ and $q_{3}$. Let $(\ell_{1},\ell_{2},\ell_{3},\ell_{3}^{\prime},\ell_{2}^{\prime},\ell_{1}^{\prime})$ be the intersections of $L$ with the zeroes of $q_{1}$, $q_{2}$ and $q_{3}$ respectively, counted with multiplicity and cyclically ordered. We fix an affine chart for $L$ such that this tuple corresponds to the tuple $x_{3}^{\prime}<x_{2}^{\prime}<x_{1}^{\prime}\leq x_{1}<x_{2}<x_{3}$ of real numbers. This yields the following :
$[\ell_{1},\ell_{3},\ell_{3}^{\prime},\ell_{1}^{\prime}]=\frac{x_{3}-x_{1}^{\prime}}{x_{3}-x_{3}^{\prime}}\times\frac{x_{1}-x_{3}^{\prime}}{x_{1}-x_{1}^{\prime}}.$
$[\ell_{1},\ell_{3},\ell_{3}^{\prime},\ell_{1}^{\prime}]=\left(\frac{x_{3}-x_{1}^{\prime}}{x_{3}-x_{2}^{\prime}}\times\frac{x_{1}-x_{2}^{\prime}}{x_{1}-x_{1}^{\prime}}\right)\times\left(\frac{x_{3}-x_{2}^{\prime}}{x_{3}-x_{3}^{\prime}}\times\frac{x_{1}-x_{3}^{\prime}}{x_{1}-x_{2}^{\prime}}\right).$
Moreover one has:
@@
$\text{cr}([q_{3}],[q_{2}])\leq[\ell_{2},\ell_{3},\ell_{3}^{\prime},\ell_{2}^{\prime}]=\frac{x_{3}-x_{2}^{\prime}}{x_{3}-x_{3}^{\prime}}\times\frac{x_{3}-x_{2}^{\prime}}{x_{3}-x_{2}^{\prime}}\leq\frac{x_{3}-x_{2}^{\prime}}{x_{3}-x_{3}^{\prime}}\times\frac{x_{1}-x_{3}^{\prime}}{x_{1}-x_{2}^{\prime}}.$
Hence one has $[\ell_{1},\ell_{3},\ell_{3}^{\prime},\ell_{1}^{\prime}]\geq\text{cr}([q_{2}],[q_{1}])\text{cr}([q_{3}],[q_{2}])$ for every such projective line $L$.
@@
A sequence of quadrics such that the cross ratio distance between the first and last quadric goes to +$\infty$+ ∞ satisfies that the intersection of all half-spaces determined by the quadrics is a projective subspace.
@@
Let $(q_{n})_{n\in\mathbb{N}}$ be a sequence of quadrics such that $q_{n+1}-q_{n}$ is positive for all $n\in\mathbb{N}$ and $\text{cr}([q_{n}],[q_{0}])$ goes to +$\infty$+ ∞. Then $\bigcap_{n\in\mathbb{N}}\{q_{n}\leq 0\}\subset\mathbb{P}(V)$ is a projective subspace.

### v2

@@ 3.1. Pencils of quadrics
For any $\Gamma$-invariant Riemannian metric $g$ on $\mathbb{S}\mathcal{E}$, any map $\sigma$ as the proof of Proposition 4.5 is a quasi-isometric embedding. Indeed let $\delta$ be the maximum of $d_{g}(x,x^{\prime})$ or $d_{H}(\sigma(x),\sigma(x^{\prime}))$ for any $x,x^{\prime}$ in the same fiber of $\operatorname{pr}:\mathbb{S}\mathcal{E}\to\widetilde{N}$. For every $x,y\in\mathbb{S}\mathcal{E}$ one can find $x^{\prime},y^{\prime}$ in the same fibers respectively as $x,y$ and in the same flow line for $\Phi$. Hence for the constants $C,D$ from Proposition 4.5:
$\frac{1}{C}d_{g}(x,y)-D-(\frac{2}{C}+2)\delta\leq d_{H}(\sigma(x),\sigma(y))\leq Cd_{g}(x,y)+D+(2C+2)\delta.$
@@ 4.2. Existence of fitting flows
Recall that $\operatorname{pr}:\mathcal{E}\to\operatorname{Gr}_{d}(\mathcal{Q})$ is the bundle map and $\pi:\mathcal{E}\to\mathcal{Q}$ is the tautological map.
@@
This lemma is an inverse of point $(iii)$ of Proposition 3.12: here we fix an element $q\in u(x)$ whereas before we were fixing a $v\in TM$. In order to construct this inverse we will use the fact that a continuous odd map between spheres of equal dimensions must have odd degree [Hat02, Proposition 2B.6] and therefore it must be surjective, as any non-surjective map into a sphere is contractible.
@@
Finally we will make this construction so that in addition $\mathrm{d}\operatorname{pr}\left(\psi(v)\right)=v$ for $v\in T_{x}M\setminus\{0\}$, but this property will be only used during the construction.
@@
[…]
If we can construct such continuous maps, the fact that $\overline{\phi}$ is an odd map between two spheres of the same dimension implies that it is homotopically non-trivial and therefore surjective. In particular for all $q]\in\mathbb{S}\mathcal{E}_{u(x)}\simeq\mathbb{S}u(x)$ there exist a $v\in T_{x}M$ such that $\phi(v)=q$. The element $\mathrm{w}=\psi(v)$ then satisfies the required conditions, so this finishes the proof.
@@
Since $u$ is a fitting immersion, the point $(iii)$ of Proposition 3.12 implies that given $v_{0}\in T_{x}M$ one can construct $\phi(v_{0})\in u^{*}\mathcal{E}_{x}$ and $\psi(v_{0})\in T_{\phi(v_{0})}u^{*}\mathcal{E}$ such that $\mathrm{d}\pi(\psi(v_{0}))$ is positive and $\mathrm{d}\operatorname{pr}\left(\psi(v_{0})\right)=v_{0}$.
Note that the condition that $\mathrm{d}\pi(\psi(v))$ is positive is an open condition and the condition that $\mathrm{d}\operatorname{pr}\left(\psi(v)\right)=v$ requires that $\psi$ is a section of an affine sub-bundle. Hence for every $v\in T_{x}M\setminus\{0\}$ we can find a small neighborhood $\mathcal{S}$ in a sphere in $T_{x}M$ containing $v_{0}$ on which we can define $\phi$ and a lift $\psi$ such that $\mathrm{d}\pi(\psi(v))$ is positive and $\mathrm{d}\operatorname{pr}\left(\psi(v)\right)=v$ for all $v\in\mathcal{S}$. We take $\mathcal{S}$ small enough so that it does not contain any antipodal pair of points.
We define $U$ to be the set of non-zero elements $\lambda v$ for all $\lambda\in\mathbb{R}$ and $v\in\mathcal{S}$, and we extend $\phi$ and $\psi$ to $U$ in a homogeneous way. We define $\phi$ on $U$ so that $\phi_{i}(\lambda v)=\lambda\phi_{i}(v)$ for all $\lambda\in\mathbb{R}$ non-zero and $v\in\mathcal{S}$. We set $\psi(\lambda v)=\lambda(\mathrm{d}m_{\lambda})\psi_{i}(v)$ for $\lambda\in\mathbb{R}$, where $m_{\lambda}$ is the multiplication by $\lambda$ on $u^{*}\mathcal{E}_{x}$. Note that $\mathrm{d}\pi\left(\psi_{i}(\lambda v)\right)=\lambda^{2}\mathrm{d}\pi\left(\psi_{i}(v)\right)$ is positive and $\mathrm{d}\operatorname{pr}\left(\psi_{i}(\lambda v)\right)=\lambda v$ for all $\lambda v\in U$. Indeed $\pi\circ m_{\lambda}=\lambda\pi$ so $\mathrm{d}\pi\circ\mathrm{d}m_{\lambda}=\lambda\mathrm{d}\pi$ and $p\circ m_{\lambda}=p$ so $\mathrm{d}\operatorname{pr}\circ\mathrm{d}m_{\lambda}=\mathrm{d}\operatorname{pr}$.
We therefore can construct an open cover $\{U_{i}\}_{i\in I}$ of $T_{x}M$ and continuous maps $\phi_{i}:U_{i}\to\mathbb{S}u^{*}\mathcal{E}\subset u^{*}\mathcal{E}$, and lifts $\psi_{i}:U_{i}\to Tu^{*}\mathcal{E}$ such that $\mathrm{d}\pi(\psi_{i}(v))$ is positive and $\mathrm{d}\operatorname{pr}\left(\psi_{i}(v)\right)=v$ for any $i\in I$ and $v\in U_{i}$. The $U_{i}$ can be assumed invariant by scalar multiplication, and $\phi$ and $\psi$ satisfy the aforementioned homogeneity conditions.
@@
Let us check that $\phi(v)$ is always non-zero. This is where we use that $\mathrm{d}\operatorname{pr}\left(\psi_{i}(v)\right)=v$ for $i\in I$, and we also use that the fitting immersion $u$ is defining a smooth fibration of the cone $S^{2}V^{>0}$. Let $\gamma:\mathbb{R}\to M$ be a curve such that $\gamma(0)=x$ and $\gamma^{\prime}(0)=v$. For all $t\in\mathbb{R}$ the intersection $u^{\circ}(\gamma(t))\cap S^{2}V^{>0}$ is a non-empty convex set, so we can construct a section $s:t\in\mathbb{R}\mapsto S^{2}V^{>0}$ of the fibration, i.e. such that for all $t\in\mathbb{R}$, $s(t)\in u^{\circ}(\gamma(t))$.
Let us fix $i\in I$. Let $q:\mathbb{R}\to\mathbb{S}u^{*}\mathcal{E}$ be such that $q^{\prime}(0)=\psi(v)$, which implies that $p\circ q(t)$ is equal to $\gamma(t)$ at the first order around $t=0$ since $\gamma^{\prime}(0)=\mathrm{d}\operatorname{pr}\left(\psi(v)\right)=v$. Since $s(t)\in u^{\circ}(\gamma(t))$, one has $\pi(q(t))\cdot s(t)=0$ at the first order around $t=0$. Taking the first derivative at $t=0$ of this equation we get:
@@
$\mathrm{d}\operatorname{pr}(\psi(v))=\sum_{i\in I}\chi_{i}(v)\mathrm{d}\operatorname{pr}(\psi_{i}(v))=\left(\sum_{i\in I}\chi_{i}(v)\right)v=v.$
@@
To prove Proposition 4.8 we use the directions $\mathrm{w}$ from Lemma 4.9 and we glue these vectors into a vector field using one again a partition of the unit. We construct the vector field $W:\mathbb{S}u^{*}\mathcal{E}\to T\mathbb{S}u^{*}\mathcal{E}$ in a similar manner as $\psi:\mathbb{S}TM\to T\mathbb{S}u^{*}\mathcal{E}$. Morally $W=\psi\circ\phi^{-1}$", but the map $\phi$ constructed previously is not a priori bijective.
@@
We construct a continuous vector field $\mathrm{W}$ over $u^{*}\mathcal{E}$, except the zero section that is homogeneous, i.e. such that for all $x\in M$ and non zero $q\in u(x)$ one has $W_{\lambda q}=\mathrm{d}(m_{\lambda})W_{q}$ for all $\lambda\in\mathbb{R}^{>0}$, where $m_{\lambda}$ is the multiplication by $\lambda$ on $\mathcal{E}$. Such a vector field defines a vector field $\overline{W}$ on $\mathbb{S}u^{*}\mathcal{E}$. We require moreover that $\mathrm{d}\pi(W)$ is always positive.
Given any non-zero $q_{0}\in u^{*}\mathcal{E}$, Lemma 4.9 provides the existence of some $\mathrm{w}_{0}\in T_{(u(x),q)}\mathcal{E}$ for all $(x,q)\in u^{*}\mathcal{E}$ such that:
  1. $(i)$
    $\mathrm{d}\operatorname{pr}(\mathrm{w}_{0})\in\mathrm{d}u(T_{x}M)$,
  2. $(ii)$
    $\mathrm{d}\pi(\mathrm{w}_{0})$ is positive.
Property $(i)$ implies that $\mathrm{w}_{0}$ defines a vector in $T_{x,q_{0}}u^{*}\mathcal{E}$. For each such $q_{0}$ one can find a neighborhood $U$ of it in $u^{*}\mathcal{E}$ that is invariant by the $\mathbb{R}^{>0}$-action, and on which one can define a map $W$ satisfying $(i)$ as well as the homogeneity condition $W_{\lambda q}=\mathrm{d}(m_{\lambda})W_{q}$. Condition $(ii)$ being an open condition invariant by the action of $m_{\lambda}$ it is satisfied automatically for $U$ small enough.
Using a partition of the unit as in Lemma 4.9 we construct the desired vector field on $u^{*}\mathcal{E}$. Note that both property $(i)$ and the homogeneity are preserved by linear combinations, and $(ii)$ is preserved by positive combinations.
@@ 5.1. The Anosov property
If we fix an affine chart of $L$ so that $\ell_{1},\ell_{2},\ell_{2}^{\prime},\ell_{1}^{\prime}$ correspond to the real numbers $x_{1}^{\prime}<x_{2}^{\prime}\leq x_{2}<x_{1}$, the cross ratio $\ell_{1},\ell_{2},\ell_{2}^{\prime},\ell_{1}^{\prime}]$ is defined as:
$\ell_{1},\ell_{2},\ell_{2}^{\prime},\ell_{1}^{\prime}]=\frac{x_{2}-x_{1}^{\prime}}{x_{2}-x_{2}^{\prime}}\times\frac{x_{1}-x_{2}^{\prime}}{x_{1}-x_{1}^{\prime}}.$ (2)
An other equivalent definition, if we take non-zero representatives $v_{1},v_{2},v_{2}^{\prime},v_{1}^{\prime}$ of $\ell_{1},\ell_{2},\ell_{2}^{\prime},\ell_{1}^{\prime}$:
$\ell_{1},\ell_{2},\ell_{2}^{\prime},\ell_{1}^{\prime}]=\frac{v_{2}\wedge v_{1}^{\prime}}{v_{2}\wedge v_{2}^{\prime}}\times\frac{v_{1}\wedge v_{2}^{\prime}}{v_{1}\wedge v_{1}^{\prime}}.$
This expression makes sense as $\bigwedge^{2}L\simeq\mathbb{R}$, and it is independent of the chosen identification. The equivalence betweens these formulations is due to the fact that if $v_{1}=e_{1}+x_{1}e_{2}$ and $v_{2}=e_{1}+x_{2}e_{2}$ then $v_{2}\wedge v_{1}=(x_{2}-x_{1})e_{2}\wedge e_{1}$.
Definition/Proposition 5.1.
Let $q_{1}],[q_{2}]\in\mathbb{S}\mathcal{Q}^{\operatorname{mix}}$ be such that for some choice of representatives, the difference $q_{2}-q_{1}$ is positive. For every projective line $L$ intersecting $\{q_{1}>0\}$ and $\{q_{2}<0\}$ we can define $\ell_{1}$ and $\ell_{1}^{\prime}$ are the zeros of $q_{1}$ on $L$ and $\ell^{\prime}_{1}$ and $\ell_{2}^{\prime}$ and the zeros of $q_{2}$ on $L$ so that the points $\ell_{1},\ell_{2},\ell_{2}^{\prime},\ell_{1}^{\prime}\in\mathbb{P}(V)$ are cyclically in this order on $L$. We define the cross-ratio distance $\text{cr}([q_{2}],[q_{1}])\in[1,\infty)$ between $q_{2}]$ and $q_{1}]$ as the minimum of the cross ratio $\ell_{1},\ell_{2},\ell_{2}^{\prime},\ell_{1}^{\prime}]$ for any such projective line $L$, which is well defined and in $(1,\infty)$.
We now verify that this minimum exists and is in $(1,\infty)$.
Proof.
First note that the assumption $q_{2}-q_{1}$ is positive implies that $\{q_{1}\geq 0\}\subset\{q_{2}>0\}$ and $\{q_{2}\leq 0\}\subset\{q_{1}<0\}$. If $L$ intersects $\{q_{1}>0\}$ and $\{q_{1}<0\}$ then the quadratic forms $q_{1}$ and $q_{2}$ are both positive on some points and negative on some points of $L$, and hence they both admit exactly two zeroes on $L$. These zeroes, denoted respectively $\ell_{1}$$\ell_{1}^{\prime}$ for $q_{1}$ and $\ell_{2}$$\ell_{2}^{\prime}$ for $q_{2}$ are in the following cyclic order on L up to exchanging $\ell_{1}$ and $\ell_{1}^{\prime}$: $(\ell_{1},\ell_{2},\ell_{2}^{\prime},\ell_{1}^{\prime})$. One can refer to Figure 1 where $\{q_{i}>0\}$ is the interior of the corresponding ellipse.
Consider the compact space $X$ of projective lines $L$ intersecting $\{q_{1}\geq 0\}$ and $\{q_{2}\leq 0\}$. For every such line we can define the lines $\ell_{1},\ell_{1}^{\prime}$ and $\ell_{2},\ell_{2}^{\prime}$ as in previously where $\ell_{1},\ell_{1}^{\prime}\neq\ell_{2},\ell_{2}^{\prime}$ but eventually $\ell_{1}=\ell_{1}^{\prime}$ or $\ell_{2}=\ell_{2}^{\prime}$ which occur respectively exactly if $L$ intersects trivially $\{q_{1}>0\}$ or $\{q_{2}>0\}$. Indeed in these cases the quadratic forms $q_{i}$ are semi-positive or semi negative on $L$ and hence have a double zero. By looking at the formula (2) on an affine chart such that $x_{1}^{\prime}<x_{2}^{\prime}\leq x_{2}<x_{1}$ we see the cross ratio $\ell_{1},\ell_{2},\ell_{2}^{\prime},\ell_{1}^{\prime}]$ is in $(1,\infty]$. This ratio is finite if and only if $\ell_{1}\neq\ell_{1}^{\prime}$ and $\ell_{2}\neq\ell_{2}^{\prime}$ which occurs exactly when $L$ belongs to the interior of $X$, i.e. when $L$ intersects $\{q_{1}>0\}$ and $\{q_{2}<0\}$. Finally note that this ratio depends continuously on $L\in X$ and that the interior of $X$ is non-empty since $q_{1}$ and $q_{2}$ are mixed. Therefore the minimum on $X$ is well defined in $(1,\infty)$ and is reached on the interior of $X$. ∎
@@
Let $q_{1}],[q_{2}],[q_{3}]\in\mathbb{S}\mathcal{Q}^{\operatorname{mix}}$ be such that $q_{3}-q_{2}$ and $q_{2}-q_{1}$ are positive. Then :
@@
Pick any projective line that crosses $\{q_{1}>0\}$ and $\{q_{3}<0\}$. Let the intersections of $L$ with the zeroes of $q_{1}$, $q_{2}$ and $q_{3}$ be respectively $(\ell_{1},\ell_{2},\ell_{3},\ell_{3}^{\prime},\ell_{2}^{\prime},\ell_{1}^{\prime})$, counted with multiplicity and cyclically ordered. We fix an affine chart for $L$ such that this tuple corresponds to the tuple $x_{3}^{\prime}<x_{2}^{\prime}<x_{1}^{\prime}\leq x_{1}<x_{2}<x_{3}$ of real numbers. This yields the following :
$\ell_{1},\ell_{3},\ell_{3}^{\prime},\ell_{1}^{\prime}]=\frac{x_{3}-x_{1}^{\prime}}{x_{3}-x_{3}^{\prime}}\times\frac{x_{1}-x_{3}^{\prime}}{x_{1}-x_{1}^{\prime}}.$
$\ell_{1},\ell_{3},\ell_{3}^{\prime},\ell_{1}^{\prime}]=\left(\frac{x_{3}-x_{1}^{\prime}}{x_{3}-x_{2}^{\prime}}\times\frac{x_{1}-x_{2}^{\prime}}{x_{1}-x_{1}^{\prime}}\right)\times\left(\frac{x_{3}-x_{2}^{\prime}}{x_{3}-x_{3}^{\prime}}\times\frac{x_{1}-x_{3}^{\prime}}{x_{1}-x_{2}^{\prime}}\right).$
Note that if $B>A>0$, for all $c>0$ one has $\frac{A}{B}\leq\frac{A+c}{B+c}$ and respectively $\frac{B+c}{A+c}\leq\frac{B}{A}$. Hence, due to the ordering of the tuple, we have the following inequalities by setting respectively $A=x_{2}-x_{1}^{\prime}$, $B=x_{2}-x_{2}^{\prime}$, $c=x_{3}-x_{2}$ and $A=x_{1}-x_{2}^{\prime}$, $B=x_{1}-x_{3}^{\prime}$, $c=x_{2}-x_{1}$:
$\frac{x_{2}-x_{1}^{\prime}}{x_{2}-x_{2}^{\prime}}\leq\frac{x_{3}-x_{1}^{\prime}}{x_{3}-x_{2}^{\prime}}\;,\;\frac{x_{2}-x_{3}^{\prime}}{x_{2}-x_{2}^{\prime}}\leq\frac{x_{1}-x_{3}^{\prime}}{x_{1}-x_{2}^{\prime}}.$
Hence we obtain:
@@
$\text{cr}([q_{3}],[q_{2}])\leq[\ell_{2},\ell_{3},\ell_{3}^{\prime},\ell_{2}^{\prime}]=\frac{x_{3}-x_{2}^{\prime}}{x_{3}-x_{3}^{\prime}}\times\frac{x_{2}-x_{3}^{\prime}}{x_{2}-x_{2}^{\prime}}\leq\frac{x_{3}-x_{2}^{\prime}}{x_{3}-x_{3}^{\prime}}\times\frac{x_{1}-x_{3}^{\prime}}{x_{1}-x_{2}^{\prime}}.$
Hence one has $\ell_{1},\ell_{3},\ell_{3}^{\prime},\ell_{1}^{\prime}]\geq\text{cr}([q_{2}],[q_{1}])\text{cr}([q_{3}],[q_{2}])$ for every such projective line $L$.
@@
A sequence of quadrics such that the cross ratio distance between the first and last quadric goes to $\infty$ satisfies that the intersection of all half-spaces determined by the quadrics is a projective subspace.
@@
Let $(q_{n})_{n\in\mathbb{N}}$ be a sequence of quadrics such that $q_{n+1}-q_{n}$ is positive for all $n\in\mathbb{N}$ and $\text{cr}([q_{n}],[q_{0}])$ goes to $\infty$. Then $\bigcap_{n\in\mathbb{N}}\{q_{n}\leq 0\}\subset\mathbb{P}(V)$ is a projective subspace.